OpenStudy (anonymous):

Guys, can you give me spare time helping me answer this Thanks, medal will be GIVEN. Solve for the unknown By Substitution 1. 5x-3y-16=0 4x-5y-5=0 By Add a Subt 1a. x-3y=2 2x+3y=13 2b. 2/x+3/y=7 6/x-6/y=1

terenzreignz (terenzreignz):

To solve for substitution, first, do you know how to isolate a variable?

terenzreignz (terenzreignz):

To solve *by* substitution... sorry.

OpenStudy (anonymous):

is this correct 5x+3y+16?

terenzreignz (terenzreignz):

I thought it was 5x - 3y - 16 = 0 ?

OpenStudy (anonymous):

yes and i change the sign, isnt that the 1st step? to find x?

terenzreignz (terenzreignz):

Sure... find x. Manipulate the equation so that x stands alone on one side, with the rest on the other.

OpenStudy (anonymous):

this lesson is linear equation with two variable btw

terenzreignz (terenzreignz):

Yup... I got that :)

OpenStudy (anonymous):

so 5x=3y+16

terenzreignz (terenzreignz):

Here, I'll give you an example... 2x - 12y + 10 = 0 So, bring everything without "x" to the right side, like so... 2x - 12y = -10 2x = 12y - 10 then divide everything by 2... x = 6y - 5 And that's how you isolate a variable :)

terenzreignz (terenzreignz):

Oh, it seems you have it down... great :D

terenzreignz (terenzreignz):

Just one more step to isolate that x.

OpenStudy (anonymous):

divide?

terenzreignz (terenzreignz):

Divide, yes, by what? :)

OpenStudy (anonymous):

can you help me on that :D

terenzreignz (terenzreignz):

?? You have \[\Large 5x = 3y + 16\] so dividing both sides by 5 should give you \[\Large x = \frac{3y+16}{5}\]

OpenStudy (anonymous):

ok i get that then is this right ?\[2\left(\begin{matrix}3y+16 \\ 5\end{matrix}\right)\]

terenzreignz (terenzreignz):

Why 2? (also, if you wanna put fractions, type \frac{numerator}{denominator} in your equation editor)

OpenStudy (anonymous):

i mean 4. sorry

terenzreignz (terenzreignz):

That's right :) \[\Large 4\left(\frac{3\color{red}y+16}{5}\right)-5\color{red}y - 5 = 0\] As you can see, now, it is just a linear equation in one variable. Solve for y, and you're good-to-go :)

OpenStudy (anonymous):

i dont know how to solve \[4\frac{3y+16}{5}\]

terenzreignz (terenzreignz):

I see... you get rid of the denominator by multiplying everything by 5, first... Or you could rely on this fact first... \[\Large p\cdot \frac{a}b = \frac{pa}b\] IE, just mulitply 4 to the numerator first.

OpenStudy (anonymous):

so it will be 4(15y+80)?

terenzreignz (terenzreignz):

Yes... techincally BUT we didn't multiply 5 just so you could distribute it to the numerator, we multiply 5 to CANCEL THE DENOMINATOR Don't confuse those two methods I mentioned... they're not the same... either you distribute the 4 first OR you multiply everything by 5. Your call.

OpenStudy (anonymous):

so my answer is valid? 4(15y+80)?

terenzreignz (terenzreignz):

that's not the answer yet, that's just a step. It is a valid step, but also unnecessary, and actually rather unhelpful, because you have to reverse what you did anyway... What I meant was this \[\Large 4(5)\left(\frac{3\color{red}y+16}{5}\right)-5(5)\color{red}y - 5(5) = 0\] So that we can do this \[\Large 4(\cancel5)\left(\frac{3\color{red}y+16}{\cancel5}\right)-25\color{red}y -25 = 0\] And so we're left with this \[\Large 4(3\color{red}y+16)-25\color{red}y - 25 = 0\]

OpenStudy (anonymous):

ok then after that we multiply now? 12y+64-25y-25?

OpenStudy (anonymous):

is this correct also? x+2y-5=0 4x+y-6=0 x=2y+5 4x+y-6=0 4(-2y+5)+y-6=0 -7 =-20+6 -7y=-14 7 7 (x=2) x=-2(2)+5 =-4+5 (y=1)

terenzreignz (terenzreignz):

Hang on a sec...

terenzreignz (terenzreignz):

It's wrong.

OpenStudy (anonymous):

why

terenzreignz (terenzreignz):

From x + 2y - 5 = 0 you got x = 2y + 5 faulty... redo that step.

terenzreignz (terenzreignz):

On second thought, that may have been a typo, as I see you used the correct substitution when it mattered... recalibrating :D

terenzreignz (terenzreignz):

It's correct. Great job :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but if i apply that method in what were working on will i be correct?

terenzreignz (terenzreignz):

Of course... you just need a refresher when it comes to dealing with fractions :)

OpenStudy (anonymous):

12y+64-25y-25? so this becomes = 37y=-89?

terenzreignz (terenzreignz):

37y? Note the minus sign next to 25y...

terenzreignz (terenzreignz):

Great... now solve for x and head-on home :D

OpenStudy (anonymous):

13y 39 13 13 = 3y

OpenStudy (anonymous):

5x-3(3)-16 -9-16? is this the right equation i think its wrong

OpenStudy (anonymous):

can you help me out with the 2 number please. specially the fraction one

terenzreignz (terenzreignz):

lol that was a pretty nice trick question, number 2... let's write it out...

terenzreignz (terenzreignz):

\[\Large \frac2x + \frac3y=7\]\[\Large \frac6x - \frac6y = 1\]

terenzreignz (terenzreignz):

This does look confusing, doesn't it? :D

OpenStudy (anonymous):

yes, but im not done finding the x in number one :D

terenzreignz (terenzreignz):

Oh, by all means... do so :D

OpenStudy (anonymous):

i dont know the equation if im right. 5x-3(3)-16 -9-16?

terenzreignz (terenzreignz):

Yeah, that's right... and all that's equal to zero... 5x - 9 - 16 = 0

OpenStudy (anonymous):

can it be x=0?

terenzreignz (terenzreignz):

Of course not. Solve for x... You solved for y earlier, and you solved it well... utilise those same skills to solve for x now :D

OpenStudy (anonymous):

but this is my equation so whats next for this?

OpenStudy (anonymous):

5x-3(3)-16

terenzreignz (terenzreignz):

5x - 9 - 16 = 0 ? Much like earlier, of course, bring everything without x to the other side, and solve.

OpenStudy (anonymous):

5x=-9-16 5x=24? can you help me out on this and before you get going can you help me solve the fraction in no 2 1st so that i can review it later.

terenzreignz (terenzreignz):

Still wrong... 5x - 9 - 16 = 0 Redo it... :P

OpenStudy (anonymous):

5x=9+16?

terenzreignz (terenzreignz):

Okay, that's better. Now solve for x.

terenzreignz (terenzreignz):

Quickly now... time's not on my side.

OpenStudy (anonymous):

i dont know whats next can you help me out

terenzreignz (terenzreignz):

Add 9 and 16?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

so y=3 x=5

OpenStudy (anonymous):

so y=3 x=5 is the final answer in 1?

terenzreignz (terenzreignz):

Good. And you're done. As for the one involving fractions, you can substitute first, if it makes it simpler. Let \[\Large \bar x = \frac1x\]\[\Large \bar y = \frac1y\] So it then becomes... \[\LARGE 2\bar x + 3\bar y = 7\]\[\LARGE 6\bar x - 6\bar y = 1\] and now it looks more familiar... solve normally from here, and then substitute back.

terenzreignz (terenzreignz):

I need to go now... I'm sure someone else would be willing to take up this challenge :) ---------------------------------------- Terence out

OpenStudy (anonymous):

x 6 - 12x+18y = 49 x 2 - 12x-12y =2?

OpenStudy (anonymous):

terenz can you leave out the solution here and answer so i can review it later use. thans